TWO forces 5 N and 7 N are acting on a point. Which of the following cannot be their resultant force? 1 N, 3 N, 7 N, 8 N. 


1
Expert's answer
2021-04-06T13:53:49-0400

Two forces can provide the maximal resultant force if and only if they are collinear and look in the same direction. In this case in order to obtain the resultant force we should add their magnitudes:


Fmax=5N+7N=12NF_{max} = 5N + 7N = 12N

Two forces can provide the minimal resultant force if and only if they are collinear and look in the opposite direction. In this case in order to obtain the resultant force we should subtract their magnitudes (the smaller one from the bigger one):


Fmin=7N5N=2NF_{min} = 7N-5N = 2N

The magnitude of any other possible resultant force is then in this range: from 2N2N to 12N12N. SInce 1N1N is out of this range, the answer is 1N1N.


Answer. 1N.